OpenStudy (anonymous):

Solve by factoring and list only the positive solution: 2x^2 - 5x = 88

OpenStudy (ikram002p):

can u factor it ?

OpenStudy (anonymous):

Doesn't it have to equal 0?

OpenStudy (anonymous):

Would it be 2x2 - 5x - 88=0? first

OpenStudy (mathstudent55):

That's correct. The first step is to put everything on the left side equaling zero like you did. Now you need to factor it.

OpenStudy (anonymous):

or you can also do middle term breaking.

OpenStudy (anonymous):

Can you explain more please?

OpenStudy (anonymous):

How would I factor 2x^2 - 5x - 88=0??

OpenStudy (anonymous):

Find the solution of the equal - find discriminant D=5^2-2*4*(-88)=.. can you find roots if you know discriminant? Roots are 8 and -11/2. so factoring will give you (x-8)*(x+11/2)=0

OpenStudy (mathstudent55):

@Keridavis You need to factor this polynomial equation: \(2x^2 - 5x - 88=0\) This is a polynomial of the type \(ax^2 + bx + c\) , where \(a \ne 1\). This is how you factor this type of polynomial: 1. Multiply ac together. In your case, a = 2, and c = -88, so ac = -176. 2. Find two numbers that multiply to ac and add to b. In your case, ac = -176, and b = -5 Since 176 = 2 * 2 * 2 * 2 * 11, we can see that -16 * 11 = -176, and -16 + 11 = -5, so our two numbers are -16 and 11. 3. Break up the middle term of the polynomial using the two numbers found in the step above. \(2x^2 - 5x - 88=0\) becomes: \(2x^2 - 16x + 11x - 88 =0\) 4. Factor by grouping. That means factor out a common factor from the first two terms, and factor out a common factor from the last two terms. \(2x^2 - 16x + 11x - 88 =0\) after factoring by grouping we get: \(2x(x - 8) + 11(x - 8)=0\) Notice we have the term (x - 8) in common, so we can factor it out leaving: \((x - 8)(2x + 11) = 0\) Now that the trinomial is factored, you can continue to solve for x by letting each factor equal zero. \(x - 8 = 0\) or \(2x + 11 = 0\) \(x = 8\) or \(2x = -11\) \(x = 8\) or \(x = -\dfrac{11}{2} \) Since you are asked to only show positive solutions, your answer is x = 8.

OpenStudy (mathstudent55):

Do you understand the method?